# Back to Bayesics: Uncovering Human Mobility Distributions and Anomalies with an Integrated Statistical and Neural Framework

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: International Conference on Advances in Geographic Information Systems 2024; October 29–November 1, 2024; Atlanta, GAisbn: 978-1-4503-XXXX-X/18/06submissionid: 216ccs: Information systems Location based servicesccs: Computing methodologies Mixture modelsccs: Computing methodologies Bayesian network modelsccs: Computing methodologies Neural networks
@@ 2.1. Anomaly Detection
Anomaly detection in spatiotemporal data has garnered significant attention, particularly for identifying irregular patterns in user behavior. Existing approaches to anomaly detection in mobility data can be broadly categorized into clustering-based, distance-based, reconstruction-based, prediction-based, and density estimation-based methods. Clustering-based methods group data into clusters and flag anomalies as outliers  (Li et al., 2021; Pu et al., 2020; Kiss et al., 2014). These methods often assume that normal data forms distinct clusters, an assumption that may not hold true in complex mobility data. Distance-based methods detect anomalies by measuring deviations from the norm  (Arias et al., 2023; Sarafijanovic-Djukic and Davis, 2019; Huo et al., 2019). While intuitive and straightforward to implement, these approaches heavily rely on the choice of distance metric — a challenge in complex datasets where different features may have varying scales or significance. Moreover, they suffer from the curse of dimensionality, where the effectiveness of distance metrics diminishes as the number of dimensions increases, making it difficult to distinguish between normal and anomalous points. Reconstruction-based methods use models like deep autoencoders to reconstruct data and identify anomalies based on reconstruction errors. Although these methods have shown promise  (Zhou and Paffenroth, 2017; Liu et al., 2020), they are prone to overfitting, where the model may learn to perfectly reconstruct the training data but fail to generalize to new, unseen data. Additionally, they are sensitive to outliers in the training data, which can lead the model to incorrectly learn to reconstruct these anomalies as normal patterns, thereby diminishing its effectiveness in identifying true anomalies.
@@ 5.1.1. Datasets
To evaluate the performance of our proposed approach, we conducted experiments using two synthetic mobility datasets, NUMOSIM-LA (Stanford et al., 2024) and Urban Anomalies (Amiri et al., 2024) that simulate human movement patterns in urban environments.
@@
Urban Anomalies [^2] The Urban Anomalies dataset can be accessed at: https://osf.io/dg6t3/.: The Urban Anomalies dataset (Amiri et al., 2024) includes synthetic simulations of urban environments in Atlanta and Berlin, each containing mobility sequences for 1000 agents, some of whom are affected by a simulated disease that causes them to eat more frequently, reduce social interactions, and stop going to work. The anomaly prevalence rates are 12% at the agent level and 12.3% at the staypoint level. Each mobility sequence includes one month of normal activities followed by a period of variable length with injected anomalies. The anomalies fall into four categories: i) Hunger: Agents visit restaurants more frequently, ii) Social: Agents visit random locations instead of meeting friends, iii) Work: Agents stop going to work, and iv) Combined: A mix of all the above anomalies. The dataset provides comprehensive information, including agent trajectories, staypoints, and social links, offering a rich set of features for evaluating anomaly detection methods.
@@ 5.1.3. Baseline
Gaussian Mixture Variational Sequence AutoEncoder (GM-VSAE) (Liu et al., 2020): A VAE-based model designed to detect trajectory anomalies through trajectory generation. The latitudes and longitudes were converted into grid indices, and were used as the only input feature to the model. This model is limited to detecting anomalies at the agent level, without providing insights into anomalies at the stay-point level.
@@ 5.1.4. Metrics
These standard metrics allow us to assess the model’s performance across all possible decision thresholds, providing a more comprehensive evaluation of anomaly detection, especially in highly imbalanced datasets where fixed thresholds may not generalize well.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 1st ACM SIGSPATIAL International Workshop on Geospatial Anomaly Detection; October 29, 2024; Atlanta, GA, USAisbn: 978-1-4503-XXXX-X/18/06submissionid: 216ccs: Information systems Location based servicesccs: Computing methodologies Mixture modelsccs: Computing methodologies Bayesian network modelsccs: Computing methodologies Neural networks
@@ 2.1. Anomaly Detection
Anomaly detection in spatiotemporal data has garnered significant attention, particularly for identifying irregular patterns in user behavior. Existing approaches to anomaly detection in mobility data can be broadly categorized into clustering-based, distance-based, reconstruction-based, prediction-based, and density estimation-based methods. Clustering-based methods group data into clusters and flag anomalies as outliers  (Li et al., 2021; Pu et al., 2020; Kiss et al., 2014). These methods often assume that normal data forms distinct clusters, an assumption that may not hold true in complex mobility data. Distance-based methods detect anomalies by measuring deviations from the norm  (Arias et al., 2023; Sarafijanovic-Djukic and Davis, 2019; Huo et al., 2019; Liu et al., 2024). While intuitive and straightforward to implement, these approaches heavily rely on the choice of distance metric — a challenge in complex datasets where different features may have varying scales or significance. Moreover, they suffer from the curse of dimensionality, where the effectiveness of distance metrics diminishes as the number of dimensions increases, making it difficult to distinguish between normal and anomalous points. Reconstruction-based methods use models like deep autoencoders to reconstruct data and identify anomalies based on reconstruction errors. Although these methods have shown promise  (Zhou and Paffenroth, 2017; Liu et al., 2020), they are prone to overfitting, where the model may learn to perfectly reconstruct the training data but fail to generalize to new, unseen data. Additionally, they are sensitive to outliers in the training data, which can lead the model to incorrectly learn to reconstruct these anomalies as normal patterns, thereby diminishing its effectiveness in identifying true anomalies.
@@ 5.1.1. Datasets
To evaluate the performance of our proposed approach, we conducted experiments using two synthetic mobility datasets, NUMOSIM-LA (Stanford et al., 2024) and Urban Anomalies (Amiri et al., 2024b, a) that simulate human movement patterns in urban environments.
@@
Urban Anomalies [^2] The Urban Anomalies dataset can be accessed at: https://osf.io/dg6t3/.: The Urban Anomalies dataset (Amiri et al., 2024b, a) includes synthetic simulations of urban environments in Atlanta and Berlin, each containing mobility sequences for 1000 agents, some of whom are affected by a simulated disease that causes them to eat more frequently, reduce social interactions, and stop going to work. The anomaly prevalence rates are 12% at the agent level and 12.3% at the staypoint level. Each mobility sequence includes one month of normal activities followed by a period of variable length with injected anomalies. The anomalies fall into four categories: i) Hunger: Agents visit restaurants more frequently, ii) Social: Agents visit random locations instead of meeting friends, iii) Work: Agents stop going to work, and iv) Combined: A mix of all the above anomalies. The dataset provides comprehensive information, including agent trajectories, staypoints, and social links, offering a rich set of features for evaluating anomaly detection methods.
@@ 5.1.3. Baseline
Gaussian Mixture Variational Sequence AutoEncoder (GM-VSAE) (Liu et al., 2020): A VAE-based model designed to detect trajectory anomalies through trajectory generation. The latitudes and longitudes were converted into grid indices, and were used as the only input to the model. This model is limited to detecting anomalies at the agent level, without providing insights into anomalies at the stay-point level.
@@ 5.1.4. Metrics
These standard metrics allow us to assess the model’s performance across all possible decision thresholds, providing a comprehensive evaluation of anomaly detection, especially in highly imbalanced datasets where fixed thresholds may not generalize well.